# vid-TLDR: Training Free Token merging for Light-weight Video Transformer

### v1

@@ Abstract
Video Transformers have become the prevalent solution for various video downstream tasks with superior expressive power and flexibility. However, these video transformers suffer from heavy computational costs induced by the massive number of tokens across the entire video frames, which has been the major barrier to training the model. Further, the patches irrelevant to the main contents, e.g., backgrounds, degrade the generalization performance of models. To tackle these issues, we propose training free token merging for lightweight video Transformer (vid-TLDR) that aims to enhance the efficiency of video Transformers by merging the background tokens without additional training. For vid-TLDR, we introduce a novel approach to capture the salient regions in videos only with the attention map. Further, we introduce the saliency-aware token merging strategy by dropping the background tokens and sharpening the object scores. Our experiments show that vid-TLDR significantly mitigates the computational complexity of video Transformers while achieving competitive performance compared to the base model without vid-TLDR. Code is available at https://github.com/mlvlab/vid-TLDR.
Figure 1: Comparison of vid-TLDR with UMT-B [33]. Without any additional training, vid-TLDR obtains comparable or even better performance (top) while reducing the considerable computational cost (bottom).
@@ 1 Introduction
With the success of Transformers in computer vision, e.g., classification [14, 52], object detection [10, 61, 43, 77, 32, 75] segmentation [59, 64], a line of works [16, 51, 57, 33, 60, 76] have proposed video Transformers to comprehend the video for various downstream tasks. The attention mechanism in Transformers shows the desirable characteristics for video understanding such as the ability to capture the spatial and temporal dependencies at the same time. Consequently, these video Transformers have been the primary backbones for the various downstream tasks in the video domain, including action recognition [73, 65], video-text retrieval [38, 17], video question-answering [63, 18], etc. Meanwhile, the self-attention scheme entails the dot-product calculation between tokens, which brings the quadratic cost in the number of tokens. This is quite problematic for previous video Transformers like UMT [33] that tokenize the whole video into a large number of tokens.
@@
In the image domain, several works have tried to mitigate the heavy computation of attention by refining the attention itself [58, 25, 6, 66], or limiting the range of attention by a pre-defined window [39, 13]. Yet, these works are not a favorable solution for video Transformers. For instance, the architecture must be altered with proposed approaches, which incurs further re-training with large datasets. Instead, recent works have proposed token reduction methods based on the flexibility of Transformers with the number of input tokens. For instance, prior works [35, 47] simply prune or merge the uninformative tokens to reduce the computational cost based on attentiveness. However, due to the low reliability of the attention in the early layer as discussed in [35], they mainly reduce the tokens in the middle layer of the Transformers. Further, we have observed that video Transformers contain a temporal bias, which makes it difficult to adopt attentiveness as the informativeness of the tokens.
@@
Based on these observations, we propose vid-TLDR, Training free token merging for Light-weight viDeo TransformeR, to effectively merge the tokens through two steps. First, we conduct saliency detection via attention sharpness. We observe that our proposed metric understands the salient region, which is more informative than backgrounds, even with the attention map in the first layer of Transformers. We also introduce the saliency-aware token merging, a training-free plug-in module to suppress the tokens irrelevant to the target tasks. Through saliency-aware token merging, we effectively drop the information of tokens in backgrounds and further contrast the informativeness of the foreground objects. Based on these components, we minimize the hindrance from the irrelevant tokens in the low layer of video Transformers. Through experiments, we show that, without any additional training, the adoption of vid-TLDR brings performance improvements of (+0.8%, +0.5%, +1.1%) with at least 39.5% lower FLOPs in UMT-B [33] on MSRVTT [68], MSVD [11], DiDeMo [2], respectively. To summarize, the contributions of vid-TLDR are presented as follows:
@@
  • Based on the saliency scores, we also propose saliency-aware token merging that performs the background drop, and foreground sharpening for adaptively adjusting the informativeness of the tokens.
@@
Figure 2: Visualization of the attention map of each method. Both attentiveness $\bar{a}$ and attention Rollout $\tilde{a}$ confused with foreground objects and the background, also have a temporal bias, resulting in overall low attention in the later frames. These problems are mitigated in our method, focusing on the object across all frames.
@@ 2 Preliminaries
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (1)
where $Q,K,V\in\mathbb{R}^{N\times C}$ are the projection of the tokens $X$ by learnable matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{C\times C}$. Given a video clip composed of $T$ frames in the resolution of $H\times W$, video Transformers first generate the tokens $X\in\mathbb{R}^{N\times C}$ by considering the clip as the set of tubes, where $N=\frac{T}{t}\times\frac{H}{P}\times\frac{W}{P}$, and ($t\times P\times P)$ is the size of each tube. All tokens in the video Transformers interact with others across the frames by spatiotemporal attention. Despite the advantage of capturing both spatial and temporal dependencies, it also demands enormous computational resources to handle the large number of tokens. Compared to one image, the computational cost for a clip is increased by $(\frac{T}{t})^{2}$ times since the cost of attention is quadratic in the number of tokens. This cost further increases as the number of frames per tube $t$ decreases.
@@
$a_{\text{cls}}=\text{softmax}\left(\frac{q_{\text{cls}}K^{\top}}{\sqrt{d}}\right),$ (2)
@@
$\bar{a}=\frac{1}{N}{\sum}^{N}_{i}\text{softmax}\left(\frac{q_{i}K^{\top}}{\sqrt{d}}\right),$ (3)
where $q_{i}$ is the query vector of $i$-th token. Further, we visualize the attention rollout [1], which is the well-known saliency detector by quantifying the flow of attention from the tokens in $l$-th layer to output, defined as
$\tilde{A}$ $={\prod}^{L}_{i=l}A^{i}\text{ and }\tilde{a}=\frac{1}{N}{\sum}^{N}_{i}\tilde{A}_{i},$ (4)
where $L$ is the number of layer in Transformers, and $A^{i}$ is the attention map in $i$-th layer. As shown in Figure 2, the attentiveness $\bar{a}$ failed to capture salient tokens due to the low reliability of the attention map in the early stage, and attention rollout $\tilde{a}$ also largely confused the foreground and background objects. Further, we have observed the temporal biases of video Transformers, e.g., the later frames tend to exhibit lower activation regardless of the importance of the frames. In this paper, we will propose a more reliable informativeness score based on the attention map that can be utilized from the earlier layers for vid-TLDR.
@@ 3 Method
Figure 3: Pipeline of vid-TLDR. (a) Given the attention map $A$, the saliency score $s$ is approximated by the sharpness function $S$ (Eq. 5). After that, we generate background drop mask $M$ (Eq. 8) to minimize the disturbance of background tokens. With $s$ and $M$, we perform foreground sharpening to generate $\hat{s}$ (Eq. 9). (b) Given the tokens $X$ and their corresponding mass $m$, we conduct the matching to group the input tokens. Following that, we update the mass $m$ to $\tilde{m}$ with $\hat{s}$ (Eq. 10) to highlight important foreground tokens and minimize the hindrance of background tokens. With updated mass $\tilde{m}$, the grouped tokens are merged into a token $m^{\prime}$ and $X^{\prime}$ (Eq. 11)
@@
In this section, we introduce vid-TLDR, Training free token merging for Light-weight viDeo transformeR. The goal of vid-TLDR is to effectively merge the tokens from the early stage by two steps: 1) Saliency detection via attention sharpness (Section 3.1), 2) Saliency-aware token merging (Section 3.2).
@@ 3.1 Saliency detection via attention sharpness
As discussed in Section 2, existing works do not reduce the tokens in the first few layers because of the low reliability. However, we believe that tokens irrelevant to the target tasks should be reduced as early as possible to minimize their adverse influence. To devise a way to identify such tokens, we first analyze the attention scores $a_{i}=\text{softmax}(\frac{q_{i}K^{\top}}{\sqrt{d}})$ in the first self-attention layer concerning the foreground and background query $q_{i}$, respectively. Figure 4 reveals that the background tokens are quite equally affected by neighboring tokens that exhibit similar patterns, whereas the tokens of foreground objects gather the information from more specific tokens showing sharper attention scores compared to backgrounds. Based on this observation, we devise a sharpness function $S$ to capture the saliency of tokens with entropy. Specifically, when the approximated saliency score of $i$-th token is defined as $s_{i}=S(a_{i})$, and the entropy of $a_{i}$ is defined as $H_{i}=-{\sum}^{N}_{j}a_{ij}\log a_{ij}$, we negate $H_{i}$ by $H^{\prime}=[-H_{1},...,-H_{N}]$ and formulate the sharpness function $S$ as
$s_{i}=S(a_{i})=\frac{H^{\prime}_{i}-\min(H^{\prime})}{\max(H^{\prime})-\min(H^{\prime})},$ (5)
Since an informative foreground token has low entropy, the saliency score $s_{i}$ of it gets higher.
Remarks. To validate that proposed saliency scores $s=S(a_{i})$ are effective for early reduction, we conduct the toy experiments by pruning the 400 tokens in each layer of UMT [33] in video-text-retrieval task with MSRVTT [68]. In Table 1, we demonstrate the advantage of early reduction compared to the base model in the first row. We believe that since the approximated saliency score $s$ detects the salient region from the early layers as shown in Figure 1(d), the tokens are reduced as early as possible to prevent the influence from the tokens irrelevant to the target tasks. As a result, when we drop the tokens in the first layer, vid-TLDR achieves the best results with the lowest FLOPs.
@@
R@10 $\uparrow$ 83.9 82.1 (--1.8) 83.8 (--0.1) 86.8 86.1 (--0.7) 86.9 ($\pm$0.0)
MSVD GFLOPs $\downarrow$ 303.3 218.7 182.2 984.6 574.5 563.1
R@1 $\uparrow$ 57.1 53.4 (--3.7) 57.6 ($+$0.5) 70.3 69.5 (--0.8) 70.5 ($+$0.2)
R@5 $\uparrow$ 81.6 77.5 (--4.1) 81.9 ($+$0.3) 89.3 88.7 (--0.6) 89.2 (--0.1)
R@10 $\uparrow$ 87.7 85.2 (--2.5) 87.9 ($+$0.2) 93.2 92.7 (--0.5) 92.8 (--0.4)
@@
DiDeMo GFLOPs $\downarrow$ 303.3 223.9 183.6 984.6 574.5 559.0
R@1 $\uparrow$ 60.6 56.1 (--4.5) 61.7 ($+$1.1) 68.1 68.0 (--0.1) 70.4 ($+$2.3)
R@5 $\uparrow$ 85.3 82.2 (--3.1) 86.1 ($+$0.8) 90.0 89.4 (--0.6) 90.5 ($+$0.5)
R@10 $\uparrow$ 91.1 90.0 (--1.1) 91.1 ($\pm$0.0) 93.4 93.8 ($+$0.4) 94.1 ($+$0.7)
@@
Table 2: Video-text retrieval on MSRVTT [68], MSVD [11], ActivityNet [9], DiDeMo [2], LSMDC [49], SSV2-Label/Template [30].
@@ 3.2 Saliency-aware token merging
Given the saliency scores $s=[s_{1},...,s_{N}]$ of the tokens, one naïve approach to reduce the tokens is simply to prune the tokens through thresholding as above. Yet, for the efficient inference of video Transformers, the tokens should be batched by pruning the same number of tokens despite the varying number of background tokens in each video. To this end, instead of simply pruning the tokens, we present saliency-aware token merging by introducing the objectiveness scores during merging.
For grouping the tokens, we leverage the bipartite soft matching in ToMe [8]. The tokens $X$ are grouped based on the bipartite soft matching between two sets $X^{\text{src}},X^{\text{dst}}\subset X$, then merges the grouped tokens into
$x^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\frac{m_{j}x_{j}}{{\sum}_{j^{\prime}}m_{j^{\prime}}},$ (6)
@@
where $\{x_{i}\}_{i\in\mathcal{G}_{i}}\subset X$, $\forall_{\mathcal{G}_{i},\mathcal{G}_{j}}\mathcal{G}_{i}\cap\mathcal{G}_{j}=\emptyset$, and $m_{i}$ is the mass of token $x_{i}$. Then, the attention is also changed with $m$ as
$A=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\log m\right).$ (7)
@@
Originally, for the mass $m$ of the tokens, ToMe use the number of constituent tokens. Although it has been proven to be effective in alleviating redundancies, it does not consider the importance of each token. Here, we redefine the mass for saliency-aware token merging based on two perspectives: 1) the hindrance imposed by the uninformative tokens should be minimized, and 2) the mass should reflect the importance of the tokens even within the foreground objects. To this end, we first minimize the hindrance from the background tokens by introducing background drop mask. Given the saliency scores, we define the mask as
$M_{i}=\begin{cases}1&\text{if }s_{i}>\bar{s},\\ 0&\text{otherwise},\end{cases}$ (8)
where $\bar{s}=\frac{\sum^{N}_{i}s_{i}}{N}$. In other words, we consider the tokens having the saliency scores lower than the average as the backgrounds and set the scores as 0. After that, with $M$, we conduct foreground sharpening to highlight more informative tokens among the foreground objects by rescaling the saliency scores as
$\hat{s}=\frac{M\odot(s-\bar{s})}{\max(M\odot(s-\bar{s}))},$ (9)
where $\odot$ indicates the element-wise multiplication. Yet, we have one more issue while adopting $\hat{s}$ as the mass. If we impose $\hat{s}=0$ to every token in the group $\mathcal{G}_{i}$, it may lose the entire information after fusion. So, we prevent this by updating the mass with $\hat{s}$ as
$\tilde{m}_{i}=\begin{cases}\hat{s}*m_{i},&\text{if }x_{i}\in X^{\text{src}}\\ m_{i}&\text{if }x_{i}\in X^{\text{dst}}\end{cases},$ (10)
Then, Equation 6 is modified as
$x^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\frac{\tilde{m}_{j}x_{j}}{{\sum}_{j^{\prime}}\tilde{m}_{j^{\prime}}}\text{ and }m^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\tilde{m}_{i},$ (11)
where $m^{\prime}_{i}$ is the mass of $i$-th fused token. Namely, as shown in Figure 3, we guarantee the information of the $X^{\text{dst}}$ while updating the mass of the source token. Note that, although the same number of tokens are merged in each video, we could further reduce the uninformative tokens based on saliency-aware token merging. We show that this background drop and foreground sharpening for token merging achieves promising improvements (see Section 4.2).
@@
UMT-B [33] 25M 51.0 44.9 58.3 61.6
UMT-L [33] 25M 58.8 58.2 66.8 70.4
UMT-B-Ours 25M 50.9 45.0 57.8 63.0
UMT-L-Ours 25M 58.1 57.8 66.5 72.4
Table 3: Text-to-video retrieval on MSRVTT (MSR.) [68], DiDeMo (DiDe.) [2], ActivityNet (Act.) [9], MSVD [11]. “#Pairs” denotes the number of pre-training pairs. Underlined results indicate the number reported in the official repository that corrects the misconfiguration of UMT [33] in MSVD.

### v2

@@ Abstract
Video Transformers have become the prevalent solution for various video downstream tasks with superior expressive power and flexibility. However, these video transformers suffer from heavy computational costs induced by the massive number of tokens across the entire video frames, which has been the major barrier to train and deploy the model. Further, the patches irrelevant to the main contents, e.g., backgrounds, degrade the generalization performance of models. To tackle these issues, we propose training-free token merging for lightweight video Transformer (vid-TLDR) that aims to enhance the efficiency of video Transformers by merging the background tokens without additional training. For vid-TLDR, we introduce a novel approach to capture the salient regions in videos only with the attention map. Further, we introduce the saliency-aware token merging strategy by dropping the background tokens and sharpening the object scores. Our experiments show that vid-TLDR significantly mitigates the computational complexity of video Transformers while achieving competitive performance compared to the base model without vid-TLDR. Code is available at https://github.com/mlvlab/vid-TLDR.
Figure 1: Comparison of vid-TLDR (Ours) with UMT [33]. Without any additional training, vid-TLDR obtains comparable or even better performance than the base model UMT (left) while reducing the considerable computational cost (right). UMT-B (87M) is used.
@@ 1 Introduction
With the success of Transformers in computer vision, e.g., classification [14, 52], object detection [10, 61, 43, 77, 32, 75], segmentation [59, 64], a line of works [16, 51, 57, 33, 60, 76] have proposed video Transformers to comprehend the video for various downstream tasks. The attention mechanism in Transformers shows the desirable characteristics for video understanding such as the ability to capture the spatial and temporal dependencies at the same time. Consequently, these video Transformers have been the primary backbones for the various downstream tasks in the video domain, including action recognition [73, 65], video-text retrieval [38, 17], video question-answering [63, 18], etc. Meanwhile, the self-attention mechanism entails the dot-product calculation between tokens, which brings the quadratic cost in the number of tokens. This poses a challenge for existing video Transformers like UMT [33] that tokenize the whole video into a large number of tokens.
@@
In the image domain, several works have tried to mitigate the heavy computation of attention by refining the attention itself [58, 25, 6, 66], or limiting the range of attention by a pre-defined window [39, 13]. Yet, these works are not a favorable solution for video Transformers since the methods entail architectural changes, requiring re-training video models with large datasets. As an alternative, in the image domain several works have proposed ‘training-free’ token reduction methods using the flexibility of Transformers in handling a variable number of input tokens. For instance, prior works [35, 47] simply prune or merge the uninformative tokens to reduce the computational cost based on attentiveness. However, we observed that the existing training-free token reduction methods for images are suboptimal for video transformers. First, previous attention-based informative scores are not accurate enough to use in early layers as discussed in [35]. So, the token reduction cannot be performed at earlier layers. Further, the attention scores of video Transformers contain a temporal bias, which makes it difficult to directly adopt them as the informativeness of the tokens, see Figure 2.
@@
Based on these observations, we propose vid-TLDR, Training-free token merging for Light-weight viDeo TransformeR, to effectively merge the tokens through two steps. First, we conduct saliency detection via attention sharpness. We observe that our proposed metric understands the salient region, which is more informative than backgrounds, even with the attention map in the first layer of Transformers. We also introduce the saliency-aware token merging, a training-free plug-in module to suppress the tokens irrelevant to the target tasks. Through saliency-aware token merging, we effectively drop the information of tokens in backgrounds and further contrast the informativeness of the foreground objects. Based on these components, we minimize the hindrance by irrelevant tokens from the early layers of video Transformers. Through experiments, we show that, without any additional training, the adoption of vid-TLDR brings performance improvements of (+0.8%, +0.5%, +1.1%) with at least 39.5% lower FLOPs in UMT-B [33] on MSRVTT [68], MSVD [11], DiDeMo [2], respectively. To summarize, the contributions of vid-TLDR are presented as follows:
@@
  • Based on the saliency scores, we also propose saliency-aware token merging with the masked saliency scores for adaptively adjusting the informativeness of the tokens.
@@
Figure 2: Visualization of the attention map of each method in the first layer. Both attentiveness $\bar{a}$ and attention Rollout $\tilde{a}$ confused with foreground objects and the background, also have a temporal bias, resulting in overall low attention in the later frames. These problems are mitigated in our method, focusing on the object across all frames.
@@ 2 Preliminaries
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{C}}\right)V,$ (1)
where $Q,K,V\in\mathbb{R}^{N\times C}$ are the projection of the tokens $X\in\mathbb{R}^{N\times C}$ by learnable matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{C\times C}$. Given a video clip composed of $T$ frames in the resolution of $H\times W$, video Transformers first generate the tokens $X\in\mathbb{R}^{N\times C}$ by considering the clip as the set of tubes, where $N=\frac{T}{t}\times\frac{H}{P}\times\frac{W}{P}$, and ($t\times P\times P)$ is the size of each tube. All tokens in the video Transformers interact with others across the frames by spatio-temporal attention. Despite the advantage of capturing both spatial and temporal dependencies, it also demands enormous computational resources to handle the large number of tokens. Compared to one image, the computational cost for a clip is increased by $(\frac{T}{t})^{2}$ times since the cost of attention is quadratic in the number of tokens. This cost further increases as the number of frames per tube $t$ decreases.
@@
$a_{\text{cls}}=\text{softmax}\left(\frac{q_{\text{cls}}K^{\top}}{\sqrt{C}}\right),$ (2)
@@
$\bar{a}=\frac{1}{N}{\sum}^{N}_{i}A_{i},$ (3)
where $A_{i}$ is the $i$-th row vector of the attention matrix $A=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{C}}\right)$. Further, we visualize the attention rollout [1], which is the well-known saliency detector by quantifying the flow of attention from the tokens in the $l$-th layer to output, defined as
$\tilde{A}^{l}$ $={\prod^{\curvearrowleft}}^{L}_{i=l}A^{i}\text{ and }\tilde{a}^{l}=\frac{1}{N}{\sum}^{N}_{i}\tilde{A}^{l}_{i},$ (4)
where $L$ is the number of layer in Transformers, and $A^{i}\in\mathbb{R}^{N\times N}$ is the attention map in $i$-th layer. As shown in Figure 2, the attentiveness $\bar{a}$ failed to capture salient tokens due to the low reliability of the attention map in the early stage, and attention rollout $\tilde{a}$ also largely confused the foreground objects. Further, we have observed the temporal biases of video Transformers, e.g., the later frames exhibit lower activation regardless of the importance of the frames.
@@ 3 Method
Figure 3: Pipeline of vid-TLDR. (a) Given the attention map $A$, the saliency score $s$ is approximated by the sharpness function $S$ (Eq. 5). After that, we generate background drop mask $M$ (Eq. 8) to minimize the disturbance of background tokens. With $s$ and $M$, we generate masked saliency scores $\hat{s}$ (Eq. 9). (b) Given the tokens $X$ and their corresponding mass $m$, we conduct the matching to group the input tokens. Following that, we update the mass $m$ to $\tilde{m}$ with $\hat{s}$ (Eq. 10) to highlight important foreground tokens and minimize the hindrance of background tokens. With updated mass $\tilde{m}$, the grouped tokens are merged into a token $m^{\prime}$ and $X^{\prime}$ (Eq. 11)
@@
We introduce vid-TLDR, Training free token merging for Light-weight viDeo transformeR. The goal of vid-TLDR is to effectively merge the tokens from the early stage by two steps: 1) Saliency detection via attention sharpness (Section 3.1), 2) Saliency-aware token merging (Section 3.2).
@@ 3.1 Saliency detection via attention sharpness
As discussed in Section 2, existing works do not reduce the tokens in the first few layers because of the low reliability. However, we believe that tokens irrelevant to the target tasks should be reduced as early as possible to minimize their adverse influence. To this end, we analyze the attention scores $A_{i}=\text{softmax}(\frac{q_{i}K^{\top}}{\sqrt{C}})$ in the first self-attention layer concerning the foreground and background tokens. Figure 4 reveals that the background tokens are quite equally affected by neighboring tokens, whereas the tokens of foreground objects gather the information from more specific tokens showing sharper attention scores compared to backgrounds. Based on this observation, we devise a sharpness function $S$ to capture the saliency of tokens with entropy. Specifically, using the negative entropy given as $H_{i}={\sum}^{N}_{j}A_{ij}\log A_{ij}$, we define the sharpness function $S$ as
$s_{i}=S(H_{i})=\frac{H_{i}-\min(H)}{\max(H)-\min(H)},$ (5)
where $H=[H_{1},\ldots,H_{N}]$. Since an informative foreground token has low entropy, its saliency score $s_{i}$ is usually high.
Remarks. To validate the saliency scores $s$, we conduct the experiments by pruning the 400 tokens in each layer of UMT [33] on video-text-retrieval task with MSRVTT [68]. In Table 1, we demonstrate that the pruning by the proposed score accelerates the base model (first row) with the competitive performance across layers. Interestingly, the token reduction in the earlier layers is more effective. To be specific, the token reduction in the first layer shows the best results (50.4%) with the lowest FLOPs (237.6 (G)). Based on this observation, we mainly applied our token reduction to earlier layers. Qualitative results in Figure 2(d) show that the saliency scores in the first layer successfully detect the salient region.
@@
R@10 $\uparrow$ 83.9 82.1 (--1.8) 83.8 (--0.1) 86.8 86.1 (--0.7) 86.9 ($+$0.1)
MSVD GFLOPs $\downarrow$ 303.3 218.7 181.3 984.6 574.5 563.1
R@1 $\uparrow$ 62.1 59.6 (--2.5) 62.7 ($+$0.6) 70.3 69.5 (--0.8) 70.4 ($+$0.1)
R@5 $\uparrow$ 84.7 83.8 (--0.9) 84.8 ($+$0.1) 89.3 88.7 (--0.6) 90.5 ($+$1.2)
R@10 $\uparrow$ 90.0 89.0 (--1.0) 89.8 (--0.2) 93.2 92.7 (--0.5) 94.0 ($+$0.8)
@@
DiDeMo GFLOPs $\downarrow$ 303.3 241.4 212.8 984.6 574.5 559.0
R@1 $\uparrow$ 62.1 57.3 (--4.8) 62.4 ($+$0.3) 70.8 68.0 (--2.8) 70.4 (--0.4)
R@5 $\uparrow$ 86.8 82.6 (--4.2) 86.2 (--0.6) 90.6 89.4 (--1.2) 90.5 (--0.1)
R@10 $\uparrow$ 92.1 89.3 (--2.8) 91.6 (--0.5) 94.5 93.8 (--0.7) 94.0 (--0.5)
@@
Table 2: Video-text retrieval on MSRVTT [68], MSVD [11], ActivityNet [9], DiDeMo [2], LSMDC [49], SSV2-Label/Template [30]. Underlined results indicate the number reported with the official repository of UMT [33] that corrects the misconfiguration of it.
@@ 3.2 Saliency-aware token merging
Given the saliency scores $s=[s_{1},...,s_{N}]$ of the tokens, our goal is to merge tokens while suppressing the influence of the irrelevant tokens with low saliency scores. We start with a brief review of a training-free token reduction method, ToMe [8]. ToMe splits tokens $X$ into two sets $X^{\text{src}},X^{\text{dst}}\subset X$ and performs the bipartite soft matching between the two sets to form token groups. For each group $\mathcal{G}_{i}$, the features are aggregated as
$x^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\frac{m_{j}x_{j}}{{\sum}_{j^{\prime}\in{\mathcal{G}_{i}}}m_{j^{\prime}}},$ (6)
@@
where $\{x_{i}\}_{i\in\mathcal{G}_{i}}\subset X$, $\forall_{\mathcal{G}_{i},\mathcal{G}_{j}}\mathcal{G}_{i}\cap\mathcal{G}_{j}=\emptyset$, and $m_{i}$ is the mass of token $x_{i}$. Then, the attention is also refined with $m$ as
$A^{\prime}_{ij}=\text{softmax}\left(A_{ij}+\log m_{j}\right),$ (7)
@@
where $A_{ij}$ is the element in $i$-th row and $j$-th column. For the mass $m$ of the tokens, ToMe uses the number of constituent tokens. Although it has been proven effective in alleviating redundancies, it cannot adaptively adjust the influence of merged tokens considering its importance. We here propose a saliency-aware token merging that estimates the mass reflecting the saliency of the tokens and then merges the tokens with their corresponding mass to minimize the hindrance induced by the uninformative tokens. First, we introduce the background drop mask to update the mass of foreground and background tokens selectively. Given the saliency scores, we define the mask as
$M_{i}=\mathbf{1}_{\{s_{i}>\bar{s}\}},$ (8)
where $\mathbf{1}$ is the indicator function and $\bar{s}=\frac{1}{N}\sum^{N}_{i}s_{i}$. Using the mask above, our framework sets the saliency scores to 0 if a token has a saliency score lower than the average saliency score. With $M$, we define masked saliency scores $\hat{s}$ to focus more on informative tokens among the foreground objects by masking and rescaling the saliency scores as
$\hat{s}_{i}=\frac{M_{i}(s_{i}-\bar{s})}{\max(M_{i}(s_{i}-\bar{s}))}.$ (9)
Yet, we have one more issue while adopting $\hat{s}$ as the mass. If all tokens in the group $\mathcal{G}_{i}$ have the saliency scores lower than $\bar{s}$, feature aggregation may result in a zero vector losing the entire information. So, to prevent this, we compute the mass with $\hat{s}$ as
$\tilde{m}_{i}=\begin{cases}\hat{s}_{i}m_{i},&\text{if }x_{i}\in X^{\text{src}}\\ m_{i}&\text{if }x_{i}\in X^{\text{dst}}\end{cases}.$ (10)
Then, feature merging of Equation 6 is modified as
$x^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\frac{\tilde{m}_{j}x_{j}}{{\sum}_{j^{\prime}\in{\mathcal{G}_{i}}}\tilde{m}_{j^{\prime}}}\text{ and }m^{\prime}_{i}={\sum}_{j\in{\mathcal{G}_{i}}}\tilde{m}_{j},$ (11)
where $m^{\prime}_{i}$ is the mass of $i$-th fused token. It is worth noting that this saliency-aware token merging can be viewed as feature aggregation only with foreground tokens since the scores of the background tokens are set to 0. In other words, it is a combination of token merging and token pruning. Although the same number of tokens are merged in each video, we dynamically adjust the influence of uninformative tokens by suppressing the mess, leading to promising improvements by the proposed saliency-aware token merging, see Section 4.2.
@@
UMT-B [33] 25M 51.0 50.8 58.3 63.7
UMT-L [33] 25M 58.8 58.2 66.8 72.5
UMT-B-Ours 25M 50.9 50.5 57.8 64.1
UMT-L-Ours 25M 58.1 57.9 66.7 72.3
Table 3: Text-to-video retrieval on MSRVTT (MSR.) [68], DiDeMo (DiDe.) [2], ActivityNet (Act.) [9], MSVD [11]. “#Pairs” denotes the number of pre-training pairs. We use the models of each dataset in Table 2 to report ours.